OpenStudy (anonymous):

Can somebody please tell me how to convert f(x) into the general, vertex form of the equation??

OpenStudy (amistre64):

if its a quadratic, then completing the square would be useful

OpenStudy (amistre64):

\[f(x)=ax^2+bx+c\] \[f(x)=a(x^2+\frac bax)+c\] \[f(x)=a(x^2+\frac bax~\pm\frac{b^2}{2a^2})+c\] \[f(x)=a(x+\frac ba)^2~-\frac{b^2a}{2a^2}+c\]

OpenStudy (amistre64):

4a^2, not 2a^2 forgot to square the half of it :/

OpenStudy (amistre64):

\[f(x)=a(x^2+\frac bax~\pm\frac{b^2}{4a^2})+c\] \[f(x)=a(x+\frac b{2a})^2~-\frac{b^2a}{4a^2}+c\] thats better ....

OpenStudy (anonymous):

Can you tell me in words? Haha. I don't understand all of that... @amistre64

OpenStudy (amistre64):

i did: "if its a quadratic, then completing the square would be useful" completing the square is a method that should be "worded out" better in your material :)

OpenStudy (anonymous):

That's what the assignment is on is completing the square haha I just don't understand how to do that. It doesn't explain it very well where I can understand it. @amistre64

OpenStudy (amistre64):

do you have an example of f(x) to play with?

OpenStudy (amistre64):

its a hard method to word out, but its pretty simple to demonstrate. With a little practice it loses all of its mystery

OpenStudy (amistre64):

if you can already find the vertex (a,b) then its just a matter of formatting the "slope" \[y = m(x-a)^2+b\] if you can determine a point in the curve, say (p,q), then its just a matter of working the setup \[q = m(p-a)^2+b\]and solving for m

OpenStudy (anonymous):

I have this: f(x) = (x - 1)(x + 1) = x^2 - 1.

OpenStudy (amistre64):

thats already in vertex form :)

OpenStudy (anonymous):

Cool! :) So... how did it get to be in the vertex form?

OpenStudy (amistre64):

you have a squared "term" and a constant .... if we take the vertex general format as:\[y=a(x-h)^2+k\] and let a=1, h=0, and k=-1 \[y=1(x-0)^2+(-1)\] \[y=x^2-1\]

OpenStudy (amistre64):

just as a sheer coincidence, the vertex form and standard form of that setup are the same

OpenStudy (anonymous):

Thanks :)